Suppose a firm’s production function is given by Q = L1/2*K1/2. The Marginal Product of Labor and the Marginal Product of Capital are given by:

MPL = ½ L-1/2K1/2and MPK = ½ L1/2K-1/2

a) Suppose the price of labor is w = 18, and the price of capital is r = 2. Derive the firm’s total cost function.

b) What is the firm’s marginal cost?

c) For this problem, you will sketch the graph of the firm’s isoquant for Q = 12 units of output, and on the same graph sketch the firm’s isocost line associated with the total cost of producing Q = 12 units of output. To get this total cost, you must use the Total Cost function from part a). Please scale your graph up to 72 units of Labor on the horizontal axis, and 72 units of Capital on the vertical axis (do not go above 72 units on either axis). For the isocost line, clearly identify the vertical and horizontal intercepts. For the isoquant, clearly identify 4 combinations of Labor and Capital that will produce Q = 12 (including the bundle that minimizes the firm’s cost of production). Make sure your graph is neatly and accurately drawn and carefully labeled.
Answer #1

A firms production function is given by ! a) o = 1/2K/2 labor . Marginal product of MPL = 20 = 2 2 ² 4/2 L2L. Marginal RoducTalla 32 Q = 1/2 k % -) Q = _Y2 (94)h a 20 a) D=L and as; K= 9L= q* 9 = 3Q Now, e substituting 2 = 0 and K= 3Q in the Cost f. Now when Q=12 Iso quart equation is 12=2%2K / 2 KM :-) When 12 = k½ = L = 4 .2=4 =) 140/2 =) 12 = K² 1 . -) K12=6 =) k= 36Now why Q = 12 Total cost ez 12 Q = (12x12) = 144 tice equation of iso cost lines : e= WLtok =) 144 = 181+2K Now horizontal i727 A 36% 14,36) 2012 n 0 : 4 8 9 72 L CS Scanned with CamScanner
Answer #2

It is solved below:

W = 18 0=2 Short vun total cost =WLtolk = 18L+QK - (2) sowing foort from equation (!-> 2 Putting this value in equation as STB. Marginal cost = dc = 12 d = 12C = 120 212.12 = 144 Total cost when 9 = 12 = Q = 12, K= 36 = 36 La H (4, 36) is one combination 36 that prodreces 12 unite SIf L=16, K = 9, 6 = 12 +2 = 4.3=12 (16,9) is the 3rd combination Similarly 9, 16) is the 4th combination Cost = WL + DIK N=18For the 3rd bundle , c = 18.16+2.9=288+18=306 For the 4th bandie, e=18.9 +2.16=162+32 = 194 so, the 1st bundle 7 (L,K) = (4,367 (4,36) (9, 1G) \(16,9) (36,4) 4 ģ 16 3 0 L intercept of the isocost line must be L=144, K=0 and K=141, L=0. only then 021